OpenStudy (anonymous):

What value of n solves the equation? 2^n = 1/8 A. –3 B. 3 C. 4 D. –4 this is 4, right?

OpenStudy (mathstudent55):

2^3 = 8 2^(-3) = 1/8

OpenStudy (anonymous):

oh.. thank you!

OpenStudy (mathstudent55):

yw